# OneSearch: A Preliminary Exploration of the Unified End-to-End Generative Framework for E-commerce Search

### v3

@@ 1. Introduction
Extensive offline evaluations are conducted on real user search logs, and the significant performance boosts demonstrate OneSearch’s effectiveness for e-commerce search. We also make multiple strict A/B testings in the KuaiShou mall search platform to verify its online effectiveness. The pure OneSearch can get comparable performance to the online MCA. By the introduction of RQ-OPQ and long behavior sequence, OneSearch can confidently improve item CTR by 1.45%, PV CTR by 1.40%. While applying the additional reward model selection can yield a 1.102% increase in search pvs, 1.67% in item CTR, 3.14% in PV CTR, 1.78% in PV CVR, 2.40% in buyers volume, and 3.22% in order volume. As a side note, we further performed an MCA testings that only included recall, pre-ranking stage, while item CTR dropped by 9.971% and order by 39.144%. These results effectively demonstrate the performance of OneSearch. Finally, the pure OneSearch can save about 75.40% operational expenditure and improve the Model Flops Utilization (MFU) from 3.26% of MCA to 27.32%. It has been successfully deployed for the entire traffic on the detail page search, 50% traffic on the mall search, and 20% traffic on the homepage search platform, for further investigation. To the best of our knowledge, it is the first industrial-deployed end-to-end generative framework for e-commerce search. We hope that exploration could further pave the way for smarter GRs in Search.
@@ 3.1.2. Core Keyword Enhancement
Item textual information often contains redundancy with many irrelevant words and even mutually exclusive attributes. While these attributes can increase item exposure, the weak semantic order caused by numerous stacked attributes makes it difficult for encoders to model key information. Here we propose using core keyword features to enhance textual representation, thereby obtaining semantic IDs dominated by these keywords. Specifically, we identified 18 structured attributes using Named Entity Recognition (NER) and mined click query-item pairs from the past 1 year as labeled data. We then compiled a list of keywords for each attribute, ranked by page views (PV) in descending order, and selected high-frequency keywords as core ones. We employ Qwen-VL (Bai et al., 2023) as the discriminant model to identify corresponding keywords for each item, while for queries, we use the Aho-Corasick Automaton (Aho and Corasick, 1975) for rapid matching during inference.
@@
This approach enhances the role of core keywords in encoding. As shown in Table 1, the core keyword enhancement scheme improves the codebook utilization rate (CUR) of RQ-Kmeans at each level and further increases the independent coding rate (ICR). For example, with a configuration of 4096-1024-512 (shown in Table 2), it results in a 0.10% CUR increment for Level 1, 24.84% for Level 2, and 26.15% for Level 3, as well as the overall ICR increasing by 6.86%.
@@ 3.2.1. Behavior Sequence Constructed User IDs
(3) $SID_{short}=\lceil\sum_{i=s_{1}}^{m}\lambda_{i}\cdot SID_{s_{i}}\rceil,\quad\text{where}\ \lambda_{i}=\frac{exp(\sqrt{i})}{\sum_{i}^{m}exp(\sqrt{i})},$
$SID_{long}=\lceil\sum_{j=l_{1}}^{n}\mu_{i}\cdot SID_{l_{i}}\rceil,\quad\text{where}\ \mu_{j}=\frac{exp(\sqrt{j})}{\sum_{j}^{n}exp(\sqrt{j})}.$
@@ 3.2.3. Implicit Long Behavior Sequence
For e-commerce platforms, user long behavior sequences primarily consist of three types: the click, the order, and the search relevant unit (RSU) (Guo et al., 2023) sequence, and the length of the behavior sequence is almost up to $10^{3}$. Therefore, it is almost impossible to integrate this information into the format of a handcrafted textual prompt. For each item within these sequences, we first map its keyword-enhanced embedding $e_{i}^{o}$ to a corresponding semantic ID (sid for simplicity), and then get RQ clustering centroid representation through the lookup method, which is considered to contain different levels of semantic information. Furthermore, we aggregate the centroids according to different levels, which not only allows the GR to systematically learn the user preferences at different levels, but also saves a lot of resources.
The overall pipeline is shown below, each item in the long-term historical sequence is replaced by the features of its RQ clustering centroid representative:
$\text{Item}_{sid}$ $=\text{RQ}(e_{i}^{o})$
$\text{Item}_{emb}$ $=\text{Emb\_lookup}(\text{Item}_{sid})$
For long-term historical sequence, overall behavior embedding is shown as
(4) $\mathbf{M}_{click}=\Bigg{\{}\sum_{i=1}^{L1}\mathbf{Item}_{emb}^{1},\sum_{i=1}^{L2}\mathbf{Item}_{emb}^{2},\sum_{i=1}^{L3}\mathbf{Item}_{emb}^{3}\Bigg{\}}$
$\mathbf{M}_{order}=\Bigg{\{}\sum_{i=1}^{L1}\mathbf{Item}_{emb}^{1},\sum_{i=1}^{L2}\mathbf{Item}_{emb}^{2},\sum_{i=1}^{L3}\mathbf{Item}_{emb}^{3}\Bigg{\}}$
$\mathbf{M}_{RSU}=\Bigg{\{}\sum_{i=1}^{L1}\mathbf{Item}_{emb}^{1},\sum_{i=1}^{L2}\mathbf{Item}_{emb}^{2},\sum_{i=1}^{L3}\mathbf{Item}_{emb}^{3}\Bigg{\}}$
$\mathbf{Q}^{(i)}=\text{QFormer}(\mathbf{M}_{click},\mathbf{M}_{order},\mathbf{M}_{RSU})$
where $\mathbf{M}_{click}$, $\mathbf{M}_{order}$, and $\mathbf{M}_{RSU}$ are referred to as click / order / RSU sequence item emb, and share the same size. To be specific, $\mathbf{M}\in\mathbb{R}^{N_{M}\times d_{model}}\ (d_{model}=768)$.
These three user behavior sequence injection methods efficiently model users’ short-term and long-term personalized preferences from different perspectives. Unlike the stacked behavior sequence concatenation used in MCA ranking stage, this modeling approach not only utilizes resources efficiently but also leverages the inference capabilities of GR models.
@@ 3.4.1. Multi-stage Supervised Fine-tuning
It should be noted that sliding window data augmentation is applied to the short behavior sequence to guide the model in learning changes in user interests and preferences. The sliding window strategy generates a new segment of the sequence and its subsequent item as the prediction target at each step by sliding a window along the user’s $Seq_{short}$ (Zhou et al., 2024). To prevent the window from becoming too large, we limit the maximum window length. This means we can augment $m$ samples for $Seq_{short}=\{s_{1},s_{2},…,s_{m}\}$, with the first sample having no sequence, and the second having the sequence with only one item $s_{1}$. This approach has been validated for achieving robust and high-performing sequential recommend and generative retrieval models (Ju et al., 2025a; Zhou et al., 2024). For e-commerce search, this also helps handle new users with limited search history by training on shorter subsequences. More details in § 4.2.
@@ 4.2. Ablation Study
Figure 7. The comparisons of MFU and OPEX for onlineMCA and OneSearch.
[…]
query suggestions.
@@ 4.4. Further Analysis
Figure 8. The online CTR relative gains for top 30 industries.
[…]
query suggestions.

### v4

@@ 1. Introduction
Extensive offline evaluations are conducted on real user search logs, and the significant performance boosts demonstrate OneSearch’s effectiveness for e-commerce search. We also make multiple strict A/B testings in the KuaiShou mall search platform to verify its online effectiveness. The pure OneSearch can get comparable performance to the online MCA. By the introduction of RQ-OPQ and long behavior sequence, OneSearch can confidently improve item CTR by 1.45%, PV CTR by 1.40%. While applying the additional reward model selection can yield a 1.10% increase in search pvs, 1.67% in item CTR, 3.14% in PV CTR, 1.78% in PV CVR, 2.40% in buyers volume, and 3.22% in order volume. As a side note, we further performed an MCA testings that only included recall, pre-ranking stage, while item CTR dropped by 9.97% and order by 39.14%. These results effectively demonstrate the performance of OneSearch. Finally, the pure OneSearch can save about 75.40% operational expenditure and improve the Model Flops Utilization (MFU) from 3.26% of MCA to 27.32%. It has been successfully deployed for the entire traffic on the detail page search, 50% traffic on the mall search, and 20% traffic on the homepage search platform, for further investigation. To the best of our knowledge, it is the first industrial-deployed end-to-end generative framework for e-commerce search. We hope that exploration could further pave the way for smarter GRs in Search.
@@ 3.1.2. Core Keyword Enhancement
Item textual information often contains redundancy with many irrelevant words and even mutually exclusive attributes. While these attributes can increase item exposure, the weak semantic order caused by numerous stacked attributes makes it difficult for encoders to model key information. Here we propose using core keyword features to enhance textual representation, thereby obtaining semantic IDs dominated by these keywords. Specifically, we identified 18 structured attributes (shown in Table 1) using Named Entity Recognition (NER) and mined click query-item pairs from the past 1 year as labeled data. We then compiled a list of keywords for each attribute, ranked by page views (PV) in descending order, and selected these high-frequency as core ones. Qwen-VL (Bai et al., 2023) is employed to identify corresponding keywords for each item, while for queries, we use the Aho-Corasick Automaton (Aho and Corasick, 1975) for rapid matching during real-time inference.
@@
This approach enhances the role of core keywords in encoding. As shown in Table 2, the core keyword enhancement scheme improves the codebook utilization rate (CUR) of RQ-Kmeans at each level and further increases the independent coding rate (ICR). For example, with a configuration of 4096-1024-512, it results in a 0.10% CUR increment for Level 1, 24.84% for Level 2, and 26.15% for Level 3, as well as the overall ICR increasing by 6.86%.
@@ 3.2.1. Behavior Sequence Constructed User IDs
(3) $SID_{short}=\lceil\sum_{i=1}^{m}\lambda_{i}\cdot SID_{s_{i}}\rceil,\quad\text{where}\ \lambda_{i}=\frac{exp(\sqrt{i})}{\sum_{i}^{m}exp(\sqrt{i})},$
$SID_{long}=\lceil\sum_{j=1}^{n}\mu_{i}\cdot SID_{l_{i}}\rceil,\quad\text{where}\ \mu_{j}=\frac{exp(\sqrt{j})}{\sum_{j}^{n}exp(\sqrt{j})}.$
@@ 3.2.3. Implicit Long Behavior Sequence
For e-commerce platforms, user long behavior sequences primarily consist of three types: the click, the order, and the search relevant unit (RSU) (Guo et al., 2023) sequence. The length of these behavior sequence is almost up to $10^{3}$. Therefore, it is almost impossible to integrate these information into the format of a handcrafted textual prompt. For each item within these sequences, we first map its keyword-enhanced embedding $e_{i}^{o}$ to a corresponding semantic ID (sid for simplicity), and then get RQ clustering centroid representation through the lookup method, which is considered to contain multiple levels of semantic information. Furthermore, we aggregate the centroids vector in the same levels, which not only allows the GR to systematically learn the user preferences at different levels, but also saves a lot of resources.
The overall pipeline is as follows, each item is represented with the aggregation of its RQ cluster centroid vectors (3 embeddings):
$\text{Item}_{sid}$ $=\text{RQ-Kmeans}(e_{i}^{o}),$
$\text{Item}_{emb}$ $=\text{Emb\_lookup}(\text{Item}_{sid}).$
For long-term historical behavior sequence, the implicit vector representation is computed as
(4) $\mathbf{M}_{click}=\Bigg\{\sum_{i=1}^{m}\mathbf{Item}_{emb}^{L_{1}},\sum_{i=1}^{m}\mathbf{Item}_{emb}^{L_{2}},\sum_{i=1}^{m}\mathbf{Item}_{emb}^{L_{3}}\Bigg\}$
$\mathbf{M}_{order}=\Bigg\{\sum_{i=1}^{n}\mathbf{Item}_{emb}^{L_{1}},\sum_{i=1}^{n}\mathbf{Item}_{emb}^{L_{2}},\sum_{i=1}^{n}\mathbf{Item}_{emb}^{L_{3}}\Bigg\}$
$\mathbf{M}_{RSU}=\Bigg\{\sum_{i=1}^{k}\mathbf{Item}_{emb}^{L1},\sum_{i=1}^{k}\mathbf{Item}_{emb}^{L_{2}},\sum_{i=1}^{k}\mathbf{Item}_{emb}^{L_{3}}\Bigg\}$
$\mathbf{Q}=\text{QFormer}(\mathbf{M}_{click},\mathbf{M}_{order},\mathbf{M}_{RSU})$
where $\mathbf{M}_{click}$, $\mathbf{M}_{order}$, and $\mathbf{M}_{RSU}$ are referred as the click / order / RSU behavior sequence vectors. $Q$ is the computed long behavior representation. To be specific, $\mathbf{Q}\in\mathbb{R}^{N_{M}\times d_{model}}\ (d_{model}=768)$.
These three user behavior sequence injection methods efficiently model users’ short-term and long-term personalized preferences from different perspectives. Unlike the stacked behavior sequence concatenation used in MCA ranking stage, this modeling approach not only utilizes resources efficiently but also fully leverages the inference capabilities of GR models.
@@ 3.4.1. Multi-stage Supervised Fine-tuning
It should be noted that sliding window data augmentation is applied to the short behavior sequence to guide the model in learning changes in user interests and preferences. The sliding window strategy generates a new segment of the sequence and its subsequent item as the prediction target at each step by sliding a window along the user’s $Seq_{short}$ (Zhou et al., 2024). To prevent the window from becoming too large, we limit the maximum window length. This means we can augment $m$ samples for $Seq_{short}=\{s_{1},s_{2},…,s_{m}\}$, with the first sample having no sequence, and the second having the sequence with only one item $s_{1}$. This approach has been validated for achieving robust and high-performing sequential recommend and generative retrieval models (Ju et al., 2025a; Zhou et al., 2024). For e-commerce search, this also helps handle new users with limited search history by training on shorter subsequences. More details are shown in § 4.2.
@@ 4.2. Ablation Study
Figure 7. The comparisons of MFU and OPEX for onlineMCA and OneSearch.:
@@ 4.4. Further Analysis
Figure 8. The online CTR relative gains for top 30 industries.: